OpenStudy (anonymous):

Simplify x^-5 by making the exponent positive. A. 1/x^5 B. 1/5x C. x5 D. 5x

geerky42 (geerky42):

What do you think?

OpenStudy (anonymous):

A

geerky42 (geerky42):

Right

OpenStudy (anonymous):

What about this? http://prntscr.com/3vjr8i @geerky42

geerky42 (geerky42):

Hint: \(\Large 64 = 4^3\)

geerky42 (geerky42):

So you have \(\Large 64^{\frac{1}{3}} \longrightarrow (4^3)^{\frac{1}{3}}\)

OpenStudy (anonymous):

D*

geerky42 (geerky42):

Exponents cancel each other, so you are left with 4. Yes it's D